Grade 12Mechanics

Three identical particles placed on corner of equilateral triangle having mass m and charge Q .if we release them simultaneously then find maximum speed of any one particle

Shiv
9 Years agoGrade 12

1 Answer

Amogh Dasture
9 Years ago
The total potential energy of the three particle system will be converted into kinetic energy at infinity which be equally divided among the three particles.
 
So,  
Total PE= \frac{3kQ^2}{r}   (As there are three particles)
 
This will be converted into total kinetic energy at infinity. Let energy of 1 particle at infinity be E
 
\therefore E=\frac{KE}{3}=\frac{3kQ^2}{3r}=\frac{kQ^2}{r}
 
Now,
 
\frac{1}{2}mv^2=\frac{kQ^2}{r}
 
\therefore v=Q \sqrt{\frac{2mk}{r}}